Calculate the acceleration of a body of mass 79 kg that has a force of 916 N acting on it at an angle of 29 degrees to the horizontal.
1
Expert's answer
2012-03-27T09:45:05-0400
The gravity force acting on a body is

G = mg = 79*9.8 = 774.2 N.

The horizontal component of a net force acting on a body is

Fh = 916*cos29 = 801.1517 N

The vertical component of a net force acting on a body is

Fv = 916*sin29 - mg = 916*sin29 - 79*9.8 = -330.1144 N.

The net force is

Fn = √(Fh²+Fv²) = √(801.1517²+330.1144²) = 8.6650 N.

So, according to the second Newton's law, the acceleration of a body is

a = Fn/m = 8.6650/79 = 0.1097 m/s².